# A General Simulation-Based Optimisation Framework for Multipoint Constant-Stress Accelerated Life Tests

### v1

@@ 6 Discussion
Estimating the lifetime of a component at its intended operating stress (design stress) is a fundamental challenge in reliability engineering. The lifetime of many modern components is such that it can be quite difficult to observe component failure in normal operating conditions. Accelerated life testing (ALT) is a tried and tested method of obtaining lifetime information in a much shorter timeframe by subjecting the components to higher-than-normal stress(es). An essential aspect of ALT is the design of the accelerated testing experiment, i.e., the accelerated test plan. Specifically, reliability practitioners are interested in obtaining the test plan configuration that maximises the lifetime information from the experiment. In other words, a key goal in ALT is to seek optimal accelerated life test plans.
In this work, we focus on test plans incorporating a single (constant) accelerating stress, and propose a simulation-based optimisation framework that utilises the differential evolution (DE) algorithm. Our general and flexible optimisation framework offers a practical solution to optimising complex multipoint designs; this is particularly useful in settings where mathematical optimisation is intractable. Using the statistical computing language R, and utilising the DE package DEoptim, we have proposed an ALT modelling/optimisation framework that can be applied across a wide range of real-world ALT testing scenarios. In some cases, it is possible to optimise the test plan by analytical means, for example, simpler test plans with a limited number of stress levels. However, theoretical optimisation becomes very challenging when one is faced with a complicated design, with multiple inputs, perhaps coupled with a complex life-stress relationship. Our simulation-based optimisation approach permits analysis of such test plans, and wherein physical constraints of the system could also be incorporated.
@@
Whilst optimisation of simpler test plans has a wide literature, discussion of complex test plans with a large number of stress levels is somewhat lacking. Our proposed framework provides insight into optimal test plan configuration irrespective of the complexity of the life-stress relationship and/or the number of design points, and hence is applicable to a variety of experimental test settings. We have demonstrated this by applying our method to lifetime data from both simulation studies (linear, quadratic, and power law life-stress curves) and real industrial life testing.
@@
Overall, our simulation-based accelerated test plan optimisation framework provides practical insights and test plans tailored to the specific application at hand. Where analytic results/guidelines exist for simpler test plans in prior literature, our results are in line with these. However, our proposal has the advantage that it can optimise any test plan, irrespective of the life-stress relationship, lifetime distribution, censoring distribution, or experimental constraints (e.g., units may need to be assigned in multiples of five or ten). This simple and effective tool is intended to supplement physical component testing and provide additional insight to the engineers and analysts working on real-life ALT problems. Incorporating this offline approach into one’s analysis may lead to a reduction in the amount of costly (in terms of both time and money) and destructive component testing, and ultimately contribute to improving component reliability.

### v2

@@ 6 Discussion
Estimating the lifetime of a component at its intended operating stress (design stress) is a fundamental challenge in reliability engineering. The lifetime of many modern components is such that it can be quite difficult to observe component failure in normal operating conditions. Accelerated life testing (ALT) is a tried and tested method of obtaining lifetime information in a much shorter time by subjecting the components to higher-than-normal stress(es). An essential aspect of ALT is the design of the accelerated testing experiment, i.e., the accelerated test plan. Specifically, reliability practitioners are interested in obtaining the test plan configuration that maximises the lifetime information from the experiment. In other words, a key goal in ALT is to seek optimal accelerated life test plans. In some cases, it is possible to optimise the test plan by analytical means, for example, simpler test plans with a limited number of stress levels. However, theoretical optimisation becomes very challenging when one is faced with a complicated design, with multiple inputs, perhaps coupled with a complex life-stress relationship.
@@
In this work, we propose a simulation-based optimisation framework to optimise ALT test plans incorporating a single (constant) accelerating stress. Whilst optimisation of simpler test plans has a wide literature, discussion of complex test plans with a large number of stress levels is somewhat lacking. Utilising the differential evolution (DE) algorithm, we discuss an optimisation framework that offers a practical solution to optimising complex multipoint designs; this is particularly useful in settings where mathematical optimisation is intractable. We have proposed a general and flexible ALT modelling/optimisation framework that, irrespective of the complexity of the test plan, can be applied across a wide range of ALT testing scenarios. We have demonstrated this by applying our method to lifetime data from both simulation studies (linear, quadratic, and power law life-stress curves) and real industrial life testing data.
The optimisation framework is intended to aid ALT practitioners in designing and optimising accelerated test plans, and as such can be easily adapted to the user’s specific testing requirements. In addition to specifying any life-stress relationship and lifetime distribution, the user can easily alter test plan parameters within the framework. For example, an analyst can modify the experiment duration, number of available test units, stress resolution, as well as incorporating experiment-specific constraints (for instance, test units may need to be assigned in specific batch/lot sizes). We consider the RMSE (root mean squared error) of out-of-sample model predictions (at a user-defined quantile) as our optimality criterion. However, the framework could be adapted to consider other optimality criteria of interest (such as the commonly-used D-optimality criterion). The simulation framework also allows one to assess a test plan’s robustness to model and/or distribution misspecification. For example, if there is uncertainty regarding model selection, an analyst could generate lifetime data adhering to a specified (true) model, and then explore the performance of models that deviate from the true functional form.
Our proposed ALT optimiser has been implemented using the statistical computing language R, and to accompany this paper, we provide open-source R code where users can utilise a range of adjustable test plan parameters to conduct ALT optimisations. In addition, the optimisation framework could be extended to incorporate additional business/experimental constraints. For example, future work could include the business cost of running experiments for extended periods, and the inherent trade-off between experiment duration and lifetime censoring. To permit users to utilise and/or adapt the optimisation tool, our code is available at: https://github.com/o-mcgrath/ALT_Optimisation
@@
Overall, our simulation-based accelerated test plan optimisation framework provides practical insights and test plan configurations tailored to the specific application at hand. Where analytic results/guidelines exist for simpler test plans in prior literature, our results are in line with these. However, our proposal has the advantage that it can optimise any test plan, irrespective of the life-stress relationship, lifetime distribution, censoring distribution, or real-world experimental constraints. This simple and effective tool is intended to supplement physical component testing and provide additional insight to the engineers and analysts working on real-life ALT problems. Incorporating this offline approach into one’s analyses may lead to a reduction of expensive (in terms of both time and money) and destructive component testing, and ultimately contribute to improving component reliability.
@@ Appendix A Empirical Models in Log-linear Form
Appendix B Further Optimisation Results (Linear AFT Model)
B.1 Increased Minimum Stress Distance - $\bm{D=10^{-1}}$
Table 5: Optimal test plans for linear AFT model. The minimum distance between consecutive stress levels is $D=10^{-1}$.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.19 0.90 6,594 7,059 169 6,794 7,337
(83) (17)
3 0.19 0.45 0.90 6,556 7,064 171 6,787 7,358
(85) (01) (14)
4 0.20 0.70 0.80 0.90 6,610 7,095 164 6,809 7,362
(83) (01) (01) (15)
5 0.20 0.41 0.70 0.80 0.90 6,511 7,197 161 6,939 7,455
(77) (02) (01) (01) (19)
6 0.20 0.31 0.60 0.70 0.80 0.90 6,651 7,236 167 6,961 7,516
(77) (05) (01) (02) (01) (14)
6* 0.20 0.90 6,752 7,179 167 6,917 7,463
(77) (23)
B.2 Increased Test Units - $\bm{n=200}$
Table 6: Optimal test plans for linear AFT model. The number of test units is $n=200$.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.20 0.90 4,556 4,950 117 4,750 5,138
(175) (25)
3 0.19 0.52 0.90 4,640 4,841 108 4,763 5,118
(172) (01) (27)
4 0.20 0.90 0.90 0.90 4,576 4,953 117 4,767 5,140
(168) (05) (25) (02)
5 0.20 0.88 0.90 0.90 0.90 4,546 4,946 109 4,764 5,116
(169) (02) (02) (02) (25)
6 0.20 0.90 0.90 0.90 0.90 0.90 4,596 5,015 115 4,827 5,205
(181) (02) (02) (01) (13) (01)
6* 0.20 0.88 0.90 0.90 4,567 4,944 114 4,750 5,134
(172) (01) (03) (24)
B.3 Reduced Simulation Replicates - $\bm{n_{\text{sim}}=500}$
Table 7: Optimal test plans for linear AFT model. The number of simulation replicates is $n_{\text{sim}}=500$.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.20 0.90 6,315 7,061 235 6,680 7,464
(84) (16)
3 0.19 0.85 0.90 6,302 7,115 233 6,741 7,502
(84) (06) (10)
4 0.20 0.89 0.90 0.90 6,286 7,073 232 6,701 7,453
(88) (07) (04) (01))
5 0.19 0.73 0.90 0.90 0.90 6,208 7,086 246 6,699 7,507
(87) (01) (02) (02) (08)
6 0.20 0.90 0.90 0.90 0.90 0.90 6,150 7,095 235 6,703 7,482
(81) (03) (02) (12) (01) (01)
6* 0.20 0.84 0.90 6,174 7,100 236 6,724 7,507
(83) (01) (16)
B.4 Reduced Test Duration - $\bm{t_{e}=6}$ months (100 test units)
Table 8: Optimal test plans for linear AFT model. The test duration is six months, and the number of test units is $n=100$.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.22 0.90 7,046 7,760 193 7,447 8,085
(80) (20)
3 0.24 0.90 0.90 7,062 7,643 177 7,360 7,936
(81) (01) (18)
4 0.23 0.75 0.90 0.90 6,961 7,711 179 7,405 7,998
(86) (01) (11) (02)
5 0.23 0.90 0.90 0.90 0.90 7,096 7,626 176 7,328 7,901
(81) (01) (02) (15) (01)
6 0.24 0.84 0.90 0.90 0.90 0.90 7,065 7,656 181 7,364 7,960
(82) (01) (01) (01) (01) (14)
6* 0.23 0.85 0.90 0.90 0.90 0.90 7,125 7,717 182 7,422 8,025
(76) (01) (01) (01) (04) (17)
B.5 Reduced Test Duration - $\bm{t_{e}=6}$ months (200 test units)
Table 9: Optimal test plans for linear AFT model. The test duration is six months, and the number of test units is $n=200$.
NN Stress Levels Min. Mean Std. $p_{0.05}$ $p_{0.95}$
$S_{1}$ $S_{2}$ $S_{3}$ $S_{4}$ $S_{5}$ $S_{6}$ RMSE RMSE Error
2 0.23 0.90 4,948 5,286 127 5,080 5,497
(169) (31)
3 0.24 0.90 0.90 4,976 5,372 123 5,172 5,579
(155) (43) (02)
4 0.23 0.66 0.90 0.90 4,919 5,308 128 5,104 5,522
(169) (01) (24) (06)
5 0.24 0.90 0.90 0.90 0.90 4,985 5,392 119 5,198 5,585
(164) (01) (28) (03) (04)
6 0.24 0.43 0.84 0.87 0.90 0.90 4,948 5,407 124 5,209 5,614
(167) (01) (01) (01) (21) (09)
6* 0.23 0.56 0.90 4,942 5,482 126 5,265 5,683
(159) (09) (32)